OpenStudy (anonymous):

Alicia bikes 25% of a 100-mi trip on the first day. She bikes 1/3 of the remaining distance on the second day. On the third day, she bikes 40% of the remaining distance. How many miles are left in Alicia’s trip?

OpenStudy (anonymous):

\[0.6 \times \Big(\frac{2}{3} \times (0.75 \times 100\text{ mi})\Big)\]

OpenStudy (anonymous):

30

OpenStudy (anonymous):

That's right :)

OpenStudy (anonymous):

yayyy